OpenStudy (linn99123):

The length of the diagonal of a square room is 13 feet. What is the approximate length of one side of the room? A. 9.2 ft B. 16.0 ft C. 18.4 ft D. 22.5 ft @ilovemathbutneedhelp

OpenStudy (anonymous):

again using the Pythagorean Theorem. We know what a square looks like, so a diagonal across it makes two triangles.

OpenStudy (linn99123):

so were spliting it in half?

OpenStudy (anonymous):

Is there a picture?

OpenStudy (linn99123):

no :/

OpenStudy (anonymous):

yah im really cunfused

OpenStudy (linn99123):

so wouldnt we just divid 13? into 2

OpenStudy (linn99123):

r 2 into 13*

OpenStudy (anonymous):

\[x^2+x^2=13^2\]

OpenStudy (anonymous):

\[2x^2=169\]

OpenStudy (linn99123):

o.0 im confused

OpenStudy (anonymous):

x = 9.19238815543 or 9.2

OpenStudy (anonymous):

 using the Pythagorean theorem again